OpenStudy (anonymous):

FIRST ANSWER GET A MEDAL The digits of a two-digit number sum to 8. When the digits are reversed, the resulting number is 18 less than the original number. What is the original number?

OpenStudy (anonymous):

I will try my hardest to help you! :)

OpenStudy (anonymous):

Let x = the 10's digit Let y = the units : 10x + y = the original two digit number : The sum of the digits of a two-digit number is 14. x + y = 14 : "If the digits are reversed, the number is 18 less than the original number." Rev no. = orig no. - 18 10y + x = 10x + y - 18 10y - y = 10x - x - 18 9y = 9x - 18 simplify, divide by 9 y = x - 2

OpenStudy (anonymous):

...

OpenStudy (anonymous):

You know...if you're going to copy and paste a way to solve a problem, least you could do is change it to match the question.

OpenStudy (anonymous):

so in other words the answer would be -2

OpenStudy (anonymous):

I did not copy and past I used my BRAIN LOL

OpenStudy (anonymous):

hope you understand!

OpenStudy (anonymous):

ok find I did

OpenStudy (anonymous):

The question states that the sum of the two-digit number is 8...not 14.

OpenStudy (anonymous):

well I tried my best and have a good day y'all :)

OpenStudy (anonymous):

You can look up abbreviations on Google. It will become your lifetime best friend.

OpenStudy (anonymous):

Okei, so let me explain how you could find the answer: You have a number, a. This number is made up of two digits, we'll use x and y. According to the problem, x + y = 8. So you need to find two numbers that, when added, give you 8. It's easier to list out the numbers that can give you this. 1 7 2 6 3 5 4 4 5 3 6 2 7 1 All these, when added with their corresponding number, equal to 8 and, when put together, they give you a two-digit number. So now, looking at the list, each set of numbers have a reversed side to it. With that, you need to see which one, when subtracted with it's reverse, gives you 18. 17 reverses to 71... 17-71=? 26 reverses to 62... 26-62=? 35 reverses to 52... 35-53=? 44 reverses to 44... 44-44=? 53 reverses to 35... 53-35=? 62 reverses to 26...62-26=? 71 reverses to 17... 71-17=? So now, try and see if you can work this out.